### v1

@@ 2 Notation and Preliminaries
This recursive system (3) suggests that the realization plan $\gamma^{i}$ is uniquely determined by the values of $\gamma^{i}(w^{i}_{I^{j}_{i}}a),(j,a)\in D_{i}$, which reflects the holistic property of the sequence form. The chance player’s realization plan $\gamma^{c}=(\gamma^{c}(w^{c}):w^{c}\in W^{c})$ is determined by $\gamma^{c}(w^{c})=\prod_{a\in w^{c}\cap A(h)}f_{c}(a|h)$, which satisfies the system (3). More notations and their descriptions can be found in Table 2. The sequence form of an extensive-form game is represented as
@@
Given a realization plan profile $\gamma=(\gamma^{i}:i\in N_{c})$, the expected payoff for player $i\in N$ at sequence $w^{i}\in W^{i}$ is defined as
$\begin{array}[]{l}g^{i}(w^{i},\gamma^{-i})=\sum\limits_{w^{-i}\in{W}^{-i}}g^{i}(w^{i},w^{-i})\prod\limits_{i_{q}\neq i}\gamma^{i_{q}}(w^{i_{q}}).\end{array}$
@@
$\begin{array}[]{l}g^{i}(\gamma)=\sum\limits_{w^{i}\in W^{i}}\gamma^{i}(w^{i})g^{i}(w^{i},\gamma^{-i}).\end{array}$
@@ 3 A Sequence-Form Characterization of Normal-Form Perfect Equilibria
This section begins by examining the relationship between mixed strategies and realization plans, laying the groundwork for the characterization of normal-form perfect equilibria. Following this, a sequence-form characterization of normal-form perfect equilibria are established by the limit of Nash equilibria within a class of perturbed games.
@@ 3.1 Relationship between Mixed Strategies and Realization Plans
von Stengel et al. [36] have conducted an preliminary exploration into the payoff equivalence between mixed strategies and realization plans. In this study, we offer a more detailed description and a rigorous proof of this relationship. Furthermore, we expand the analysis by incorporating an examination of best response strategies.
Consider an extensive-form game, $\Gamma$, with $\Gamma_{n}$ as its normal form and $\Gamma_{s}$ as its sequence form. Given any pure strategy $s^{i}\in S^{i}$ of player $i\in N_{c}$, we define $s^{i}(w^{i})=\prod_{a\in w^{i}}s^{i}(a)$ for $w^{i}\in W^{i}$. For any $\sigma\in\Xi$, let $\gamma(\sigma)=(\gamma^{i}(\sigma^{i};w^{i}):i\in N_{c},w^{i}\in W^{i})$, where $\gamma^{i}(\sigma^{i};w^{i})=\sum_{s^{i}\in S^{i}}s^{i}(w^{i})\sigma^{i}(s^{i})$. It follows that $\gamma^{i}(s^{i};w^{i})=s^{i}(w^{i})$ and $\gamma(\sigma)\in\Lambda$. This relation is captured by the set $T=\left\{(\sigma,\gamma)\left|\sigma\in\Xi,\gamma=\gamma(\sigma)\right.\right\}$, which leads to the following lemma.
@@
For $\gamma\in\text{int}(\Lambda)$, we define $\sigma(\gamma)=(\sigma^{i}(\gamma^{i};s^{i}):i\in N_{c},s^{i}\in S^{i}),\text{where }\sigma^{i}(\gamma^{i};s^{i})=\prod_{j\in M_{i},a\in A(I^{j}_{i}),s^{i}(a)=1}\gamma^{i}(w^{i}_{I^{i}_{j}}a)/\gamma^{i}(w^{i}_{I^{j}_{i}})$. Given any $\gamma^{*}\in\Lambda$, let $\{\gamma^{k}\in\text{int}(\Lambda)\}_{k=1}^{\infty}$ be a convergent sequence such that $\lim_{k\rightarrow\infty}\gamma^{k}=\gamma^{*}$. It is obvious that $\sigma(\gamma^{k})\in\text{int}(\Xi)$ and $(\sigma(\gamma^{k}),\gamma^{k})\in T$. Since $\{(\sigma(\gamma^{k}),\gamma^{k})\}_{k=1}^{\infty}$ is contained within the compact set $T$, the sequence has a convergent subsequence. Denote the mixed strategy component in the limit of this subsequence as $\sigma^{*}\in\Xi$, we have $(\sigma^{*},\gamma^{*})\in T$. This completes the proof. ∎
@@
For a mixed strategy profile $\sigma\in\Xi$, the probability of reaching each terminal history $h\in Z$ is $\prod_{i\in N_{c}}\sum_{s^{i}\in S^{i}}s^{i}(w^{i}_{h})\sigma^{i}(s^{i})$. For a realization plan profile $\gamma\in\Lambda$, it is $\prod_{i\in N_{c}}\gamma^{i}(w^{i}_{h})$. When $(\sigma,\gamma)\in T$, we have $\gamma^{i}(w^{i}_{h})=\sum_{s^{i}\in S^{i}}s^{i}(w^{i}_{h})\sigma^{i}(s^{i})$ for every player $i\in N_{c}$. As a result, the probabilities that reaching each terminal history $h\in Z$ under the strategies $x$ and $\gamma$ coincide, which implies that $u^{i}(\sigma)=u^{i}(\gamma)$. This finalizes the proof. ∎
After establishing the payoff equivalence between the two type of strategies, proceed to analyze the relationship between their best responses, which underpin the proof presented in the subsequent subsection. Given $\gamma\in\Lambda$, we define the expected payoff, leading by the sequence $w^{i}\in W^{i}$, for player $i\in N$ as
$\begin{array}[]{l}g^{i}(\gamma;w^{i})=\sum\limits_{\tilde{w}^{i}\in W^{i},w^{i}\subseteq\tilde{w}^{i}}\gamma^{i}(\tilde{w}^{i})g^{i}(\tilde{w}^{i},\gamma^{-i}).\end{array}$
@@
Consider a realization plan profile $\gamma\in\Lambda$. For any $i\in N,j\in M_{i},a\in A(I^{j}_{i})$, we refer to $w^{i}_{I^{j}_{i}}a$ as an $I^{j}_{i}$-best-response sequence to $\gamma$ if the following equality holds for any $a^{\prime}\in A(I^{j}_{i})$
$\begin{array}[]{l}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j}_{i}}a)\geq\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j}_{i}}a^{\prime}).\end{array}$
We define $w^{i}_{I^{j}_{i}}a$ as a best-response sequence to $\gamma$ if, for any $j_{q}\in M_{i},a_{q}\in A(I^{j_{q}}_{i})$ with $a_{q}\in w^{i}_{I^{j}_{i}}a$, $w^{i}_{I^{j_{q}}_{i}}a_{q}$ qualifies as an $I^{j_{q}}_{i}$-best-response sequence to $\gamma$.
@@
    For any $j\in M_{i},a\in A(I^{j}_{i})$ with $s^{i}(w^{i}_{I^{j}_{i}}a)=1$, $w^{i}_{I^{j}_{i}}a$ is a best-response sequence to $\gamma$.
@@
$\bm{(1)\Rightarrow(2)}$: Assume (1) holds and (2) does not hold, there exists $j\in M_{i},a\in A(I^{j}_{i})$ with $s^{i}(w^{i}_{I^{j}_{i}}a)=1$ satisfying that $w^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma$. This means that, for some $j_{q}\in M_{i},a_{q}\in A(I^{j_{q}}_{i})$ with $a_{q}\in w^{i}_{I^{j}_{i}}a$, $w^{i}_{I^{j_{q}}_{i}}a_{q}$ is not a $I^{j_{q}}_{i}$-best-response sequence to $\gamma$. There exists $a^{\prime}_{q}\in A(I^{j_{q}}_{i})$ such that
$\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j_{q}}_{i}}a_{q})<\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j_{q}}_{i}}a^{\prime}_{q}),$
which brings a pure strategy $\tilde{s}^{i}$ such that $\gamma^{i}(\tilde{s}^{i})\in\arg\max_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j_{q}}_{i}}a^{\prime}_{q})$ and $\gamma^{i}(\tilde{s}^{i};w^{i})=\gamma^{i}(s^{i};w^{i})$ for any $w^{i}\in W^{i}$ with $a_{q},a^{\prime}_{q}\notin w^{i}$. As a result, $g^{i}(\gamma^{i}(s^{i}),\gamma^{-i})<g^{i}(\gamma^{i}(\tilde{s}^{i}),\gamma^{-i})$, which, according to Lemma 2, implies that $u^{i}(s^{i},\sigma^{-i})<u^{i}(\tilde{s}^{i},\sigma^{-i})$, thereby resulting in a contradiction.
$\bm{(2)\Rightarrow(1)}$: Assume $(2)$ holds, for any $j\in M_{i},a\in A(I^{j}_{i})$ with $s^{i}(w^{i}_{I^{j}_{i}}a)=1$, we have
$\begin{array}[]{ll}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j}_{i}}a)&=\sum\limits_{j_{q}\in M_{i}(w^{i}_{I^{j}_{i}}a)}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}\sum\limits_{a_{q}\in A(I^{j_{q}}_{i})}\tilde{\gamma}^{i}(w^{i}_{I^{j_{q}}_{i}}a_{q})g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j_{q}}_{i}}a_{q})+g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i})\\ &=\sum\limits_{j_{q}\in M_{i}(w^{i}_{I^{j}_{i}}a)}\sum\limits_{a_{q}\in A(I^{j_{q}}_{i})}s^{i}(w^{i}_{I^{j_{q}}_{i}}a_{q})\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j_{q}}_{i}}a_{q})+g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i})\\ \end{array}$
As a result of the forward induction, we can derive that
$\begin{array}[]{ll}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i})&=\sum\limits_{j\in M_{i}(\emptyset)}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}\sum\limits_{a\in A(I^{j}_{i})}\tilde{\gamma}^{i}(w^{i}_{I^{j}_{i}}a)g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j}_{i}}a)+g^{i}(\emptyset,\gamma^{-i})\\ &=\sum\limits_{j\in M_{i}(\emptyset)}\sum\limits_{a\in A(I^{j}_{i})}s^{i}(w^{i}_{I^{j}_{i}}a)\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};w^{i}_{I^{j}_{i}}a)+g^{i}(\emptyset,\gamma^{-i})\\ &=\sum\limits_{w^{i}\in W^{i}}s^{i}(w^{i})g^{i}(w^{i},\gamma^{-i})\\ &=g^{i}(\gamma^{i}(s^{i}),\gamma^{-i}).\end{array}$
@@ 3.2 A Sequence-Form Characterization of Normal-Form Perfect Equilibria
We begin by formulating a perturbation. Let $\varepsilon>0$ be a sufficiently small parameter and define the vector $\eta(\varepsilon)=(\eta^{i}(\varepsilon;w^{i}):i\in N,w^{i}\in W^{i})$, subject to the following constraints,
$\begin{array}[]{l}\sum\limits_{a\in A(I^{j}_{i})}\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)-\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ 0<\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)\leq\varepsilon,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (4)
The existence of such an $\eta(\varepsilon)$ is guaranteed by the recursiveness. Specifically, the assignment $\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)=\varepsilon^{|w^{i}_{I^{j}_{i}}a|}$ for $i\in N,(j,a)\in D_{i}$ provides a viable solution that adheres to the conditions. Given a perturbation $\eta(\varepsilon)$ satisfying the system (4), let $\Lambda(\varepsilon)=\times_{i\in N}{\Lambda^{i}(\varepsilon)}$ represent the set of perturbed realization plan profiles defined by $\Lambda^{i}(\varepsilon)=\{\gamma^{i}(\varepsilon)|\gamma^{i}(\varepsilon)\in\Lambda^{i},\gamma^{i}(\varepsilon;w^{i})\geq\eta^{i}(\varepsilon;w^{i}),w^{i}\in W^{i}\}$. We then construct a perturbed game in sequence form, denoted by $\Gamma_{s}(\varepsilon)$, where the optimal strategy for player $i$ with the strategies of other players fixed at $\hat{\gamma}^{-i}(\varepsilon)\in\Lambda^{-i}(\varepsilon)$ is determined by solving the linear optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(\varepsilon)}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}(\varepsilon))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)\leq\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (5)
Two clarifications regarding the optimization problem (5) are necessary. Firstly, we omit the payoff associated with the empty sequence in the objective function, as it does not depend on $\gamma^{i}(\varepsilon)$. Secondly, we exclude redundant inequalities in the constraints that arise from the structural similarities between $\eta^{i}(\varepsilon)$ and $\gamma^{i}(\varepsilon)$.
@@
$\begin{array}[]{l}g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(\varepsilon))+\lambda^{i}(w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=0,\;i\in N,(j,a)\in D_{i},\\ g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(\varepsilon))-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,(j,a)\notin D_{i},\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ (\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)-\eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a))\lambda^{i}(w^{i}_{I^{j}_{i}}a)=0,\\ \eta^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a)\leq\gamma^{i}(\varepsilon;w^{i}_{I^{j}_{i}}a),\;0\leq\lambda^{i}(w^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i},\end{array}$ (6)
where $\zeta^{i}_{I^{j}_{i}}(a)=\sum_{j_{q}\in M_{i}(w^{i}_{I^{j}_{i}}a)}\nu^{i}_{I^{j_{q}}_{i}}$. Consequently, $\gamma^{*}(\varepsilon)$ is a Nash equilibrium of $\Gamma_{s}(\varepsilon)$ if and only if there exists a pair $(\lambda^{*},\nu^{*})$ alongside $\gamma^{*}(\varepsilon)$ that satisfies the system (6). Following Lemma 3, we derive a specific condition that $\gamma^{*}(\varepsilon)$ must fulfill, as stated in Lemma 4.

### v3

@@ 2 Preliminaries
This recursive system (3) suggests that the realization plan $\gamma^{i}$ is uniquely determined by the values of $\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a),(j,a)\in D_{i}$, where no information set of player $i$ succeeds $\varpi^{i}_{I^{j}_{i}}a$. It reflects the holistic property of the sequence form. The chance player’s realization plan $\gamma^{c}=(\gamma^{c}(\varpi^{c}):\varpi^{c}\in W^{c})$ is determined by $\gamma^{c}(\varpi^{c})=\prod_{a\in\varpi^{c}\cap A(h)}f_{c}(a|h)$, which satisfies the system (3). More notations and their descriptions can be found in Table 2. The sequence form of an extensive-form game is represented as
@@
Given a realization plan profile $\gamma=(\gamma^{i}:i\in N_{c})$, the expected payoff for player $i\in N$ at sequence $\varpi^{i}\in W^{i}$ is defined as
$\begin{array}[]{l}g^{i}(\varpi^{i},\gamma^{-i})=\sum\limits_{\varpi^{-i}\in{W}^{-i}}g^{i}(\varpi^{i},\varpi^{-i})\prod\limits_{i_{q}\neq i}\gamma^{i_{q}}(\varpi^{i_{q}}).\end{array}$
@@
$\begin{array}[]{l}g^{i}(\gamma)=\sum\limits_{\varpi^{i}\in W^{i}}\gamma^{i}(\varpi^{i})g^{i}(\varpi^{i},\gamma^{-i}).\end{array}$
@@ 3 A Sequence-Form Characterization of Normal-Form Perfect Equilibria
This section begins by examining the relationship between mixed strategies and realization plans, laying the groundwork for the characterization of normal-form perfect equilibria. Following this, a sequence-form characterization of normal-form perfect equilibria are established as a limit point of a sequence of Nash equilibria within a class of perturbed games.
@@ 3.1 Relationship between Mixed Strategies and Realization Plans
von Stengel et al. [VonStengelComputingNormalForm2002] have conducted an preliminary exploration into the payoff equivalence between mixed strategies and realization plans. In this study, we offer a more detailed description and a rigorous proof of this relationship. Furthermore, we expand the analysis by incorporating an examination of best response strategies.
[…]
@@
For $\gamma\in\text{int}(\Lambda)$, define $\sigma(\gamma)=(\sigma^{i}(\gamma^{i};s^{i}):i\in N_{c},s^{i}\in S^{i}),\text{where }\sigma^{i}(\gamma^{i};s^{i})=\prod_{j\in M_{i},a\in A(I^{j}_{i}),s^{i}(a)=1}\gamma^{i}(\varpi^{i}_{I^{i}_{j}}a)/\gamma^{i}(\varpi^{i}_{I^{j}_{i}})$. Consider any $\gamma^{*}\in\Lambda$, and let $\{\gamma^{k}\in\text{int}(\Lambda)\}_{k=1}^{\infty}$ be a sequence converging to $\gamma^{*}$, i.e., $\lim_{k\rightarrow\infty}\gamma^{k}=\gamma^{*}$. It follows that $\sigma(\gamma^{k})\in\text{int}(\Xi)$ and $(\sigma(\gamma^{k}),\gamma^{k})\in T$ for each $k$. Because the sequence $\{(\sigma(\gamma^{k}),\gamma^{k})\}_{k=1}^{\infty}$ lies within the compact set $T$, there exists a convergent subsequence. Denote the limit mixed strategy of this subsequence as $\sigma^{*}\in\Xi$, we have $(\sigma^{*},\gamma^{*})\in T$. This completes the proof. ∎
@@
For a mixed strategy profile $\sigma\in\Xi$, the probability of reaching each terminal history $h\in Z$ is $\prod_{i\in N_{c}}\sum_{s^{i}\in S^{i}}s^{i}(\varpi^{i}_{h})\sigma^{i}(s^{i})$. For a realization plan profile $\gamma\in\Lambda$, it is $\prod_{i\in N_{c}}\gamma^{i}(\varpi^{i}_{h})$. When $(\sigma,\gamma)\in T$, we have $\gamma^{i}(\varpi^{i}_{h})=\sum_{s^{i}\in S^{i}}s^{i}(\varpi^{i}_{h})\sigma^{i}(s^{i})$ for every player $i\in N_{c}$. As a result, the probabilities that reaching each terminal history $h\in Z$ under the strategies $x$ and $\gamma$ coincide, which implies that $u^{i}(\sigma)=u^{i}(\gamma)$. This finalizes the proof. ∎
After establishing the payoff equivalence between the two type of strategies, we proceed to analyze the relationship between their best responses, which underpin the proof presented in the subsequent subsection. Given $\gamma\in\Lambda$, we define the expected payoff, leading by the sequence $\varpi^{i}\in W^{i}$, for player $i\in N$ as
$\begin{array}[]{l}g^{i}(\gamma;\varpi^{i})=\sum\limits_{\tilde{\varpi}^{i}\in W^{i},\varpi^{i}\subseteq\tilde{\varpi}^{i}}\gamma^{i}(\tilde{\varpi}^{i})g^{i}(\tilde{\varpi}^{i},\gamma^{-i}).\end{array}$
@@
Consider a realization plan profile $\gamma\in\Lambda$. For any $i\in N,j\in M_{i},a\in A(I^{j}_{i})$, we refer to $\varpi^{i}_{I^{j}_{i}}a$ as an $I^{j}_{i}$-best-response sequence to $\gamma$ if the following equality holds for any $a^{\prime}\in A(I^{j}_{i})$,
$\begin{array}[]{l}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};\varpi^{i}_{I^{j}_{i}}a)\geq\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};\varpi^{i}_{I^{j}_{i}}a^{\prime}).\end{array}$
We define $\varpi^{i}_{I^{j}_{i}}a$ as a best-response sequence to $\gamma$ if, for any $j_{q}\in M_{i},a_{q}\in A(I^{j_{q}}_{i})$ with $a_{q}\in\varpi^{i}_{I^{j}_{i}}a$, $\varpi^{i}_{I^{j_{q}}_{i}}a_{q}$ qualifies as an $I^{j_{q}}_{i}$-best-response sequence to $\gamma$.
@@
    For any $j\in M_{i},a\in A(I^{j}_{i})$ with $s^{i}(\varpi^{i}_{I^{j}_{i}}a)=1$, $\varpi^{i}_{I^{j}_{i}}a$ is a best-response sequence to $\gamma$.
@@
$\bm{(1)\Rightarrow(2)}$: Assume (1) holds and (2) does not hold, there exists $j\in M_{i},a\in A(I^{j}_{i})$ with $s^{i}(\varpi^{i}_{I^{j}_{i}}a)=1$ satisfying that $\varpi^{i}_{I^{j}_{i}}a$ is not a best-response sequence to $\gamma$. This means that, for some $j_{q}\in M_{i},a_{q}\in A(I^{j_{q}}_{i})$ with $a_{q}\in\varpi^{i}_{I^{j}_{i}}a$, $\varpi^{i}_{I^{j_{q}}_{i}}a_{q}$ is not a $I^{j_{q}}_{i}$-best-response sequence to $\gamma$. There exists $a^{\prime}_{q}\in A(I^{j_{q}}_{i})$ such that
$\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};\varpi^{i}_{I^{j_{q}}_{i}}a_{q})<\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};\varpi^{i}_{I^{j_{q}}_{i}}a^{\prime}_{q}),$
which brings a pure strategy $\tilde{s}^{i}$ such that $\gamma^{i}(\tilde{s}^{i})\in\arg\max_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};\varpi^{i}_{I^{j_{q}}_{i}}a^{\prime}_{q})$ and $\gamma^{i}(\tilde{s}^{i};\varpi^{i})=\gamma^{i}(s^{i};\varpi^{i})$ for any $\varpi^{i}\in W^{i}$ with $a_{q},a^{\prime}_{q}\notin\varpi^{i}$. As a result, $g^{i}(\gamma^{i}(s^{i}),\gamma^{-i})<g^{i}(\gamma^{i}(\tilde{s}^{i}),\gamma^{-i})$, which, according to Lemma 2, implies that $u^{i}(s^{i},\sigma^{-i})<u^{i}(\tilde{s}^{i},\sigma^{-i})$, thereby resulting in a contradiction.
$\bm{(2)\Rightarrow(1)}$: Assume $(2)$ holds, for any $j\in M_{i},a\in A(I^{j}_{i})$ with $s^{i}(\varpi^{i}_{I^{j}_{i}}a)=1$, we have
[…]
The second equation follows directly from the condition (2). As a result of the forward induction, we can derive that
$\begin{array}[]{ll}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i})&=\sum\limits_{j\in M_{i}(\emptyset)}\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}\sum\limits_{a\in A(I^{j}_{i})}\tilde{\gamma}^{i}(\varpi^{i}_{I^{j}_{i}}a)g^{i}(\tilde{\gamma}^{i},\gamma^{-i};\varpi^{i}_{I^{j}_{i}}a)+g^{i}(\emptyset,\gamma^{-i})\\ &=\sum\limits_{j\in M_{i}(\emptyset)}\sum\limits_{a\in A(I^{j}_{i})}s^{i}(\varpi^{i}_{I^{j}_{i}}a)\max\limits_{\tilde{\gamma}^{i}\in\Lambda^{i}}g^{i}(\tilde{\gamma}^{i},\gamma^{-i};\varpi^{i}_{I^{j}_{i}}a)+g^{i}(\emptyset,\gamma^{-i})\\ &=\sum\limits_{\varpi^{i}\in W^{i}}s^{i}(\varpi^{i})g^{i}(\varpi^{i},\gamma^{-i})\\ &=g^{i}(\gamma^{i}(s^{i}),\gamma^{-i}).\end{array}$
@@ 3.2 A Sequence-Form Characterization of Normal-Form Perfect Equilibria
We begin by formulating a perturbation vector. Let $\varepsilon>0$ be a sufficiently small parameter and define a vector $\eta(\varepsilon)=(\eta^{i}(\varepsilon;\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$, subject to the following constraints,
$\begin{array}[]{l}\sum\limits_{a\in A(I^{j}_{i})}\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)-\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ 0<\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)\leq\varepsilon,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (4)
The existence of such an $\eta(\varepsilon)$ is guaranteed by the recursiveness. Specifically, the assignment $\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)=\varepsilon^{|\varpi^{i}_{I^{j}_{i}}a|}$ for $i\in N,(j,a)\in D_{i}$ provides a viable solution that adheres to the conditions. Given a perturbation vector $\eta(\varepsilon)$ satisfying the system (4), let $\Lambda(\varepsilon)=\times_{i\in N}{\Lambda^{i}(\varepsilon)}$ represent the set of perturbed realization plan profiles defined by $\Lambda^{i}(\varepsilon)=\{\gamma^{i}(\varepsilon)|\gamma^{i}(\varepsilon)\in\Lambda^{i},\gamma^{i}(\varepsilon;\varpi^{i})\geq\eta^{i}(\varepsilon;\varpi^{i}),\varpi^{i}\in W^{i}\}$. We then construct a perturbed game in sequence form, denoted by $\Gamma_{s}(\varepsilon)$, where the optimal strategy for player $i$ with the strategies of other players fixed at $\hat{\gamma}^{-i}(\varepsilon)\in\Lambda^{-i}(\varepsilon)$ is determined by solving the linear optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(\varepsilon)}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,\hat{\gamma}^{-i}(\varepsilon))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)\leq\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (5)
Two clarifications regarding the optimization problem (5) are necessary. Firstly, we omit the payoff associated with the empty sequence in the objective function, as it does not depend on $\gamma^{i}(\varepsilon)$. Secondly, we exclude redundant inequalities in the constraints that arise from the same recursive equation for $\eta^{i}(\varepsilon)$ and $\gamma^{i}(\varepsilon)$, but a smaller value of $\eta^{i}(\varepsilon)$ is used at the empty sequence.
@@
$\begin{array}[]{l}g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(\varepsilon))+\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=0,\;i\in N,(j,a)\in D_{i},\\ g^{i}(\varpi^{i}_{I^{j}_{i}}a,\gamma^{-i}(\varepsilon))-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,(j,a)\notin D_{i},\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ (\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)-\eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a))\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=0,\\ \eta^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a)\leq\gamma^{i}(\varepsilon;\varpi^{i}_{I^{j}_{i}}a),\;0\leq\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i},\end{array}$ (6)
where $\zeta^{i}_{I^{j}_{i}}(a)=\sum_{j_{q}\in M_{i}(\varpi^{i}_{I^{j}_{i}}a)}\nu^{i}_{I^{j_{q}}_{i}}$. Hence, $\gamma^{*}(\varepsilon)$ is a Nash equilibrium of $\Gamma_{s}(\varepsilon)$ if and only if there exists a corresponding pair $(\lambda^{*},\nu^{*})$ such that together with $\gamma^{*}(\varepsilon)$ they fulfill the system (6). Following Lemma 3, we derive a specific condition that $\gamma^{*}(\varepsilon)$ must fulfill, as stated in Lemma 4.